OpenStudy (anonymous):

Write the equation of the line that passes through (-4, 12) and (7, 23). A. y = x + 16 B. y = 3x + 18 C. y = 1/6 x + 16 D. y = 1/3 x + 18

OpenStudy (secret-ninja):

The formula is y=mx+b The b is the y intercept, and the m is the slope.

OpenStudy (secret-ninja):

The formula to find slope is y-1y/x-x1=m

OpenStudy (anonymous):

so it cant be a?

OpenStudy (anonymous):

whats the first step?

OpenStudy (secret-ninja):



OpenStudy (anonymous):

confused now

OpenStudy (secret-ninja):

By what?

OpenStudy (anonymous):

by that equation u showed me

OpenStudy (secret-ninja):

That equation is the formula to find the slope of a line. You use the two points.

OpenStudy (secret-ninja):

Now you have the slope, 11/11 which is just 1 and when its just one, you don't put any number for slope.

OpenStudy (anonymous):

ok

OpenStudy (secret-ninja):

So: y=x+b would count on this occasion.

OpenStudy (secret-ninja):

So it would be A.

OpenStudy (anonymous):

so how would i use that equation for this question

OpenStudy (anonymous):

ok yay thank you

OpenStudy (secret-ninja):

No problem. :]

OpenStudy (anonymous):

can i ask another?

OpenStudy (secret-ninja):

Sure, just post in a new question. :)

OpenStudy (anonymous):

ok i will if i need more help

OpenStudy (secret-ninja):

Okay. :]